# SSVQ: Unleashing the Potential of Vector Quantization with Sign-Splitting

### v1

@@ Abstract
Vector Quantization (VQ) has emerged as a prominent weight compression technique, showcasing substantially lower quantization errors than uniform quantization across diverse models, particularly in extreme compression scenarios. However, its efficacy during fine-tuning is limited by the constraint of the compression format, where weight vectors assigned to the same codeword are restricted to updates in the same direction. Consequently, many quantized weights are compelled to move in directions contrary to their local gradient information. To mitigate this issue, we introduce a novel VQ paradigm, Sign-Splitting VQ (SSVQ), which decouples the sign bit of weights from the codebook. Our approach involves extracting the sign bits of uncompressed weights and performing clustering and compression on all-positive weights. We then introduce latent variables for the sign bit and jointly optimize both the signs and the codebook. Additionally, we implement a progressive freezing strategy for the learnable sign to ensure training stability. Extensive experiments on various modern models and tasks demonstrate that SSVQ achieves a significantly superior compression-accuracy trade-off compared to conventional VQ. Furthermore, we validate our algorithm on a hardware accelerator, showing that SSVQ achieves a 3× speedup over the 8-bit compressed model by reducing memory access.
@@ 1 Introduction
Deep Neural Networks (DNNs) have demonstrated remarkable success across various computer vision tasks, including image classification, object detection, segmentation, and generation. The proliferation of edge intelligence has driven significant interest in deploying DNNs to edge devices, particularly for real-time and privacy-sensitive applications. However, the rapid growth of model parameters often exceeds the capacity of accelerator on-chip SRAMs [28, 31], necessitating frequent off-chip DRAM accesses that substantially increase latency and power consumption.
@@
Fine-tuning is always essential for extremely low-bit quantization. Although vector quantization has significantly smaller quantization errors compared to uniform quantization, its potential during fine-tuning is limited. In vector quantization, to maintain the compressed format, only the codebook can be fine-tuned rather than the weights themselves. This means that all weight vectors clustered to the same codeword can only be updated in the same direction, greatly limiting the fine-tuning potential. Numerial weight subvectors are forced to move in suboptimal places which finally leads to accuracy degradation.
@@
  • We introduce a novel VQ paradigm, Sign-Spitting VQ, to boost accuracy gains during fine-tuning.
@@ 2 Related Work
Parameter-efficient Networks. In recent years, the design of parameter-efficient neural networks has gained significant attention due to the increasing demand for models that perform well under resource constraints. CNNs have seen substantial advancements with architectures like MobileNets [29, 11, 32] and EfficientNets [35], which employ techniques such as depthwise separable convolutions and neural architecture search. InceptionNeXt [43] integrates multi-scale Inception-like modules to speed up large-kernel-baed models. On the other hand, transformers have revolutionized natural language processing and are being adapted for vision tasks. EfficientViT [19] and MobileViT [25] introduce self-attention approximations that maintain linear complexity with minor impacts on accuracy. Additionally, hybrid architectures that combine the strengths of CNNs and transformers, such as MobileFormer [4], FastViT [39], have emerged as powerful solutions.
@@
Figure 1: Overview of our Sign-Splitting Vector Quantization (SSVQ) algorithm. Firstly, the sign bit is extracted and k-means clustering is applied to all-positive weights. Secondly, the fixed sign is replaced by the latent parameter’s sign function, which is updated based on STE grad approximation. Finally, an improved iterative freezing strategy is employed on the learnable sign to prevent oscillations from hurting training stability. The final storage consists of codebooks, assignments, and sign masks, which are used to reconstruct weights.
@@ 3.1 Codebook Initialization
Let $W\in R^{o\times i}$ denote the weight matrix of a linear layer, we firstly split $W$ into subvectors $w_{i,j}\in R^{d\times 1}$:
$W=\begin{bmatrix}w_{1,1}&w_{1,2}&\cdots&w_{1,i}\\ \vdots&\vdots&\ddots&\vdots\\ w_{o/d,1}&w_{o/d,2}&\cdots&w_{o/d,i}\end{bmatrix},$ (1)
For a typical vector quantization algorithm, $C=\{c_{1},...,c_{k}\}$ denotes the codebook containing $k$ codewords of size $d$. We can view $W$ as a collection of $\frac{o\times i}{d}$ subvectors of dimension $d$. Each subvector $w_{i,j}\in W$ is mapped to one of the codewords $c_{i}\in C$. The quantized weight $W_{q}$ is obtained as follow:
$W_{q}=C[A]=\begin{bmatrix}c(a_{1,1})&c(a_{1,2})&\cdots&c(a_{1,i})\\ \vdots&\vdots&\ddots&\vdots\\ c(a_{o/d,1})&c(a_{o/d,2})&\cdots&c(a_{o/d,i})\end{bmatrix}.$ (2)
To minimize the quantization error, the most popular approach is the k-means [21, 23] clustering algorithm.
3.2 Storage analysis
The compression ratio is denoted as the key metric. Denote $b_{f}$ as the bit-width of full-precision weight. Assuming that the codebook is quantized to $q_{c}$ bits, the Compression Ratio of $W$ after employing vector quantization is calculated as:
$C.R.=\frac{o*i*b_{f}}{b_{a}+b_{c}}=\frac{d*b_{f}}{\lceil\log_{2}(k)\rceil+\frac{k\times d^{2}\times q_{c}}{o*i}}$ (3)
where $b_{a}$ and $b_{c}$ are the storage costs of assignments and codebooks, respectively. $ceil$ is the function that rounds a given value up to the nearest integer.
@@ 4.1 Training limitation in Vector Quantization
Figure 2: Comparison between uniform quantization and vector quantization on how the quantized weight is updated. We assume a simple update $w^{*}=w-G_{w}$ and $c^{*}=c-G_{c}$
We begin by analyzing the differences between uniform quantization and vector quantization in quantization-aware training. As illustrated in Fig. 2 (left), UQ maintains trainable float point weights $w$, and the gradients $G_{w_{q}}$ from quantized weights are propagated to $G_{w}$ through the Straight-Through Estimator (STE). The updated $w^{*}$ are then fake quantized to obtain the updated $w_{q}^{*}$, allowing each quantized weight to be updated according to its gradient direction.
In contrast, VQ’s compressed format restricts updates to codewords only. As shown in Fig. 2 (right), the gradient of a codeword $G_{c}$ is the sum (in pytorch implementation) of the gradients of the weight vectors clustered to it. This constraint forces all quantized weights within a cluster to share the same update direction. More critically, gradient dominance frequently occurs, where a few strong gradients dictate updates for the entire cluster - as visualized in the orange part of Fig. 2, where the bottom weight’s gradient forces the upper weights to move downward. Our empirical analysis on MobileNet-V2 (d=4, k=64) confirms this as a common phenomenon in conventional VQ.
Subset grad Top grad Bottum grad
@@
Table 1: Empirical study on MobileNet-V2 model with conventional VQ. We report the cosine similarity between different subset grad and the codeword grad.
@@
Table 2: Empirical analysis of conventional VQ on MobileNet-V2: Top-1 ImageNet accuracy after 3 training epochs. Layer annotations: BN (batch normalization layers), FC (fully connected layers), Full (codebooks + BN + FC). Results indicate limited accuracy gains from codebook training.
For each codeword, we compute two gradient statistics from its clustered subvectors: (1) the sum of the top 5-10% gradients with the largest magnitudes, and (2) the sum of the bottom 50-60% gradients. Tab. 1 presents the cosine similarity between these gradient statistics and codeword gradients, revealing distinct correlation patterns that confirm gradient dominance - where a minority of high-magnitude gradients dictate codeword updates, often forcing the majority of weights into suboptimal positions. This fundamental limitation directly impacts fine-tuning effectiveness, as evidenced by the constrained accuracy improvements brought by training codewords shown in Tab. 2.
@@ 4.2 Sign-Splitting Vector Quantization
To unleash the potential of vector quantization during the fine-tuning process, we need a new compression paradigm that enables each quantized weight to update in its own gradient direction. To address this, we propose Sign-Splitting Vector Quantization (SSVQ). First, we extract the sign bit from $W$ and then apply k-means clustering to the absolute values of the weights.
$C,A=\textbf{kmeans}(|W|,k),S=\textbf{sign}(W)$ (4)
where $S$ is a 1/-1 mask with the same size as $W$. Then the compressed can be expressed as:
$W_{q}=C[A]\circ S$ (5)
Then, we can make the sign bit learnable. We can initialize a floating-point parameter $Ls$ (learnable sign) as $Ls=W\times\theta$, where $\theta$ is a constant. In this scenario, the forward computation process operates as follows:
$W_{q}=C[A]\circ sign(Ls)=\\ \begin{bmatrix}c(a_{1,1})\circ s(Ls_{1,1})&\cdots&c(a_{1,i})\circ s(Ls_{1,i})\\ \vdots&\ddots&\vdots\\ c(a_{o/d,1})\circ s(Ls_{o/d,1})&\cdots&c(a_{o/d,i})\circ s(Ls_{o/d,i})\end{bmatrix}$ (6)
During backpropagation, we jointly learn the codebook and $Ls$. We employ the Straight Through Estimator to approximate the gradients. Specifically, the gradient of $Ls$ is computed as
$g_{Ls}=\frac{\nabla L}{\nabla W_{q}}\times\frac{\nabla W_{q}}{\nabla sign(Ls)}\times\frac{\nabla sign(Ls)}{\nabla Ls}\approx\frac{\nabla L}{\nabla W_{q}}\times c$ (7)
While the direct application of STE to approximate gradients lacks precision, the gradient of $Ls$ incorporates the magnitude of the corresponding codeword. This inclusion, to some extent, helps maintain training stability, as supported by the findings in [20].
Figure 3: Comparision between VQ and SSVQ. SSVQ enables compressed weight to update in a different direction with the codeword using learnable signs, which VQ can’t.
The final storage consists of three components: the sign mask, the indices, and the codebook. The compression ratio is calculated as:
$C.R.=\frac{o*i*b_{f}}{b_{a}+b_{c}+b_{s}}=\frac{d*b_{f}}{d+\lceil\log_{2}(k)\rceil+\frac{k\times d^{2}\times q_{c}}{o*i}}$ (8)
@@
Therefore, SSVQ enables the expansion of the quantization dimension into higher-dimensional spaces while utilizing smaller codebooks. The inherent properties of clustering effectively offset the 1-bit storage overhead introduced by the sign representation.
@@ 4.3 Enhanced Iterative Freezing of Learnable Signs
5:     Update $Ls$ using $g_{Ls}^{t}=\frac{\nabla L}{\nabla W_{q}}\times c$
6:     $s^{t}=\text{sign}(Ls^{t})\neq\text{sign}(Ls^{t-1})$
@@
9:         $p_{c}=p_{c}+(\text{sign}(Ls^{t})>0)$
10:         $n_{c}=n_{c}+(\text{sign}(Ls^{t})<0)$
@@
[…]
@@
  • As shown in Fig. 4(C), frozen result based on sign-EMA also exhibits frequent oscillations. Since the sign involves only binary values, simply counting the positive and negative signs as the criterion for freezing yields more stable results, as shown in Fig. 4(B). The oscillations in the frozen results, based on majority sign voting, are significantly reduced.
Based on these observations, we propose an enhanced iterative freezing strategy for learnable signs, as summarized in Algorithm. 1. In each iteration, we update freq-EMA $f$ and record the positive $p_{c}$ and negative counts $n_{c}$ for learnable signs that have undergone sign changes. For every $Fi$ iteration, we freeze the learnable signs whose oscillation frequency exceeds the threshold and determine the frozen result based on majority sign voting. We apply a cosine schedule to the freezing threshold similar to [27].
@@ 5 Experiments
Figure 5: Quantization results of image classification on ImageNet dataset. We present the accuracy-compression trade-off curves for both VQ and SSVQ, with competing methods annotated as discrete points for comparative analysis.

### v2

@@ Abstract
Vector Quantization (VQ) has emerged as a prominent weight compression technique, showcasing substantially lower quantization errors than uniform quantization across diverse models, particularly in extreme compression scenarios. However, its efficacy during fine-tuning is limited by the constraint of the compression format, where weight vectors assigned to the same codeword are restricted to updates in the same direction. Consequently, many quantized weights are compelled to move in directions contrary to their local gradient information. To mitigate this issue, we introduce a novel VQ paradigm, Sign-Splitting VQ (SSVQ), which decouples the sign bit of weights from the codebook. Our approach involves extracting the sign bits of uncompressed weights and performing clustering and compression on all-positive weights. We then introduce latent variables for the sign bit and jointly optimize both the signs and the codebook. Additionally, we implement a progressive freezing strategy for the learnable sign to ensure training stability. Extensive experiments on various modern models and tasks demonstrate that SSVQ achieves a significantly superior compression-accuracy trade-off compared to conventional VQ. Furthermore, we validate our algorithm on a hardware accelerator, showing that SSVQ achieves a 3$\times$ speedup over the 8-bit compressed model by reducing memory access. Our code is available at https://github.com/list0830/SSVQ.
@@ 1 Introduction
Deep Neural Networks (DNNs) have demonstrated remarkable success across various computer vision tasks, including image classification [37, 24, 40, 42], object detection [46], segmentation [2, 3], and generation [35, 13]. The proliferation of edge intelligence has driven significant interest in deploying DNNs to edge devices, particularly for real-time and privacy-sensitive applications. However, the rapid growth of model parameters often exceeds the capacity of accelerator on-chip SRAMs [32, 36], necessitating frequent off-chip DRAM accesses that substantially increase latency and power consumption.
@@
Fine-tuning is always essential for extremely low-bit quantization. Although vector quantization has significantly smaller quantization errors compared to uniform quantization, its potential during fine-tuning is limited. In vector quantization, to maintain the compressed format, only the codebook can be fine-tuned rather than the weights themselves. This means that all weight vectors clustered to the same codeword can only be updated in the same direction, greatly limiting the fine-tuning potential. Consequently, many weight subvectors are updated towards suboptimal values, which ultimately degrades model accuracy.
@@
  • We introduce a novel VQ paradigm, Sign-Splitting VQ, to boost accuracy gains during fine-tuning.
@@ 2 Related Work
Parameter-efficient Networks. In recent years, the design of parameter-efficient neural networks has gained significant attention due to the increasing demand for models that perform well under resource constraints. CNNs have seen substantial advancements with architectures like MobileNets [33, 12, 37] and EfficientNets [40], which employ techniques such as depthwise separable convolutions and neural architecture search. InceptionNeXt [49] integrates multi-scale Inception-like modules to speed up large-kernel-based models. On the other hand, transformers have revolutionized natural language processing and are being adapted for vision tasks. EfficientViT [22] and MobileViT [29] introduce self-attention approximations that maintain linear complexity with minor impacts on accuracy. Additionally, hybrid architectures that combine the strengths of CNNs and transformers, such as MobileFormer [5], FastViT [45], have emerged as powerful solutions.
@@
Figure 1: Overview of our Sign-Splitting Vector Quantization (SSVQ) algorithm. Firstly, the sign bit is extracted and k-means clustering is applied to all-positive weights. Secondly, the fixed sign is replaced by the latent parameter’s sign function, which is updated based on STE grad approximation. Finally, an improved iterative freezing strategy is employed on the learnable sign to prevent oscillations from destabilizing training. The final storage consists of codebooks, assignments, and sign masks, which are used to reconstruct weights.
@@ 3 Preliminaries
Let $W\in\mathbb{R}^{O\times I}$ denote the weight matrix of a linear layer. We first partition $W$ into $N=\frac{O\times I}{d}$ subvectors $\{\mathbf{w}_{n}\}_{n=1}^{N}$, where each subvector $\mathbf{w}_{n}\in\mathbb{R}^{d}$. For simplicity, we can reshape and view $W$ as:
$W\Leftrightarrow\{\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{N}\}.$ (1)
@@ 3.1 Codebook Initialization
In vector quantization, a codebook $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{K}\}$ is maintained, containing $K$ codewords, where each $\mathbf{c}_{k}\in\mathbb{R}^{d}$. The mapping is defined by an assignment function $A:\{1,\dots,N\}\to\{1,\dots,K\}$. The quantized weight tensor $W_{q}$ is reconstructed by replacing each subvector $\mathbf{w}_{n}$ with its assigned codeword $\mathbf{c}_{A(n)}$:
$W_{q}\Leftrightarrow\{\mathbf{c}_{A(1)},\mathbf{c}_{A(2)},\dots,\mathbf{c}_{A(N)}\}.$ (2)
To minimize the quantization error $\|W-W_{q}\|^{2}$, the standard approach is to use the k-means [25, 27] algorithm.
3.2 Codebook Finetuning
To recover model accuracy after quantization, the codebook $\mathcal{C}$ is fine-tuned. Let $L$ be the network’s loss function. For a specific codeword $\mathbf{c}_{k}\in\mathcal{C}$, let $I_{k}=\{n\mid A(n)=k\}$ be the set of indices of subvectors assigned to it. The gradient for $\mathbf{c}_{k}$ is typically approximated by accumulating the gradients of its assigned subvectors, and gradient-based learning is performed as follows :
$\mathbf{c}_{k}\leftarrow\mathbf{c}_{k}-O(\sum_{n\in I_{k}}\frac{\partial L}{\partial\mathbf{w}_{n}},\theta)$ (3)
where $O(,)$ is the optimizer with hyperparameter $\theta$.
@@ 4.1 Training limitation in Vector Quantization
Figure 2: Comparison between uniform quantization and vector quantization on how the quantized weight is updated. We assume a simple update rule, e.g., $w*\leftarrow w-G_{w}$ and $c*\leftarrow c-G_{c}$.
We begin by analyzing the differences between uniform and vector quantization in quantization-aware training. As illustrated in Fig. 2 (left), UQ maintains trainable float point weights $w$, and the gradients $G_{w_{q}}$ from quantized weights are propagated to $G_{w}$ through the Straight-Through Estimator. The updated $w^{*}$ are then fake quantized to obtain the updated $w_{q}^{*}$, allowing each quantized weight to be updated according to its gradient direction.
In contrast, VQ’s compressed format restricts updates to codewords only. As shown in Fig. 2 (right), the gradient of a codeword $G_{c}$ is the sum (in pytorch implementation) of the gradients of the weight vectors clustered to it. This forces all quantized weights within a cluster to share the same update direction. More critically, gradient dominance frequently occurs, where a few strong gradients dictate the update for the entire cluster - as visualized in the orange box of Fig. 2, where the bottom weight’s gradient forces the upper weights to move downward.
To verify this, we conduct an empirical analysis on MobileNet-V2 (with $d=4,k=64$). For each codeword, we compute two gradient statistics from its clustered subvectors: (1) the sum of the top 5-10% gradients with the largest magnitudes, and (2) the sum of the bottom 50-60% gradients. Tab. 1 presents the cosine similarity between gradient subsets and the final codeword gradient. The high correlation with top gradients and low correlation with bottom gradients confirm that a minority of high-magnitude gradients dominate the update, often forcing the majority of weights into suboptimal positions. This fundamental limitation directly impacts fine-tuning effectiveness, as evidenced by the constrained accuracy improvements brought by training codewords shown in Tab. 2.
Subset grad Top grad Bottom grad
@@
Table 1: Empirical study on MobileNet-V2 model with conventional VQ. We report the cosine similarity between different subset grad and the codeword grad, confirming gradient dominance.
@@
Table 2: Fine-tuning analysis of conventional VQ on MobileNet-V2: Top-1 ImageNet accuracy after 3 training epochs. Layer annotations: BN (batch normalization layers), FC (fully connected layers), Full (codebooks + BN + FC). Results indicate limited accuracy gains from codebook training.
@@ 4.2 Sign-Splitting Vector Quantization
[…]
$\mathcal{C},A=\operatorname{k-means}(|W|,K),\quad S=\operatorname{sign}(W)$ (4)
where $S\in\{-1,1\}^{O\times I}$ is the sign mask. The quantized weight $W_{q}$ is reconstructed by combining the clustered magnitudes with their signs:
$W_{q}=\mathcal{C}[A]\circ S$ (5)
To allow for independent update directions, we make the sign mask learnable. We introduce a continuous, learnable variable $L_{s}$ initialized as $L_{s}\leftarrow\alpha\cdot W$ where $\alpha$ is a constant. The forward pass then becomes:
$W_{q}=\mathcal{C}[A]\circ\operatorname{sign}(L_{s}).$ (6)
During backpropagation, we jointly learn the codebook $\mathcal{C}$ and the sign parameter $Ls$. We employ the Straight Through Estimator to approximate the gradients for the non-differentiable sign function. Specifically, the gradient of $Ls$ is computed as
$g_{Ls}=\frac{\nabla L}{\nabla W_{q}}\circ\frac{\nabla W_{q}}{\nabla sign(Ls)}\circ\frac{\nabla sign(Ls)}{\nabla Ls}\approx\frac{\nabla L}{\nabla W_{q}}\circ c$ (7)
While the direct application of STE to approximate gradients lacks precision, the gradient of $Ls$ incorporates the magnitude of the corresponding codeword. This inclusion, to some extent, helps stabilize training, as supported by the findings in [23].
Figure 3: Comparison between VQ and SSVQ. SSVQ enables a quantized weight to update in a different direction with the codeword using learnable signs, which VQ can’t.
@@
Therefore, SSVQ enables the expansion of the quantization dimension into higher-dimensional spaces while utilizing smaller codebooks. The inherent properties of clustering effectively offset the 1-bit storage overhead introduced by the sign representation. At the same compression ratio, SSVQ has a similar clustering error compared to VQ. A detailed analysis is provided in the Supple. Material (Sec. 9).
@@ 4.3 Enhanced Iterative Freezing of Learnable Signs
5:  Update $Ls$ using $g_{Ls}^{t}=\frac{\nabla L}{\nabla W_{q}}\circ c$
6:  $s^{t}=\operatorname{sign}(Ls^{t})\neq\operatorname{sign}(Ls^{t-1})$
@@
9:    $p_{c}=p_{c}+(\operatorname{sign}(Ls^{t})>0)$
10:    $n_{c}=n_{c}+(\operatorname{sign}(Ls^{t})<0)$
@@
When a sign flips, the value of the quantized weight changes by a magnitude of $2|\mathbf{c}|$. This large, discrete change causes the learnable signs to oscillate, as both $+|\mathbf{c}|$ and $-|\mathbf{c}|$ may be suboptimal approximations. A similar issue was observed in uniform quantization-aware training [31], where weights oscillate between quantization grid points. They proposed using two Exponential Moving Averages (EMAs) to separately track the oscillation frequency and the frozen results. Inspired by  [31], we further analyze the sign dynamics in SSVQ and identify two key phenomena:
@@
  • As shown in Fig. 4(C), the frozen result based on sign-EMA also exhibits frequent oscillations. Since the sign is binary, a simpler majority vote is more robust. Fig. 4(B) shows that basing the frozen sign on a simple majority count of positive vs. negative signs results in a much more stable outcome.
Based on these observations, we propose an enhanced iterative freezing strategy for learnable signs, as summarized in Algorithm 1. In each iteration, we update freq-EMA $f$ and record the positive $p_{c}$ and negative counts $n_{c}$ for learnable signs that have undergone sign changes. For every $Fi$ iteration, we identify signs whose oscillation frequency exceeds a decaying threshold and freeze them permanently to their majority-voted sign. We apply a cosine schedule to the freezing threshold, similar to [31].
@@ 5 Experiments
Figure 5: Quantization results of classification on ImageNet dataset. We present the accuracy-compression trade-off curves for both VQ and SSVQ, with competing methods annotated as discrete points for comparative analysis. The full-precision results are marked as ’baseline’.
@@
5.1 Storage Analysis
The compression ratio (CR) is denoted as the key metric. Denote $b_{f}$ as the bit-width of full-precision weight. Assuming that the codebook is quantized to $q_{c}$ bits, the CR of $W\in\mathbb{R}^{O\times I}$ after employing conventional VQ is calculated as:
$\text{CR}_{\text{VQ}}=\frac{O\cdot I\cdot b_{f}}{b_{a}+b_{c}}=\frac{d_{1}\cdot b_{f}}{\lceil\log_{2}(k_{1})\rceil+\frac{k_{1}\cdot d_{1}^{2}\cdot q_{c}}{O\cdot I}}$ (8)
where $b_{a}$ and $b_{c}$ are the storage costs of assignments and codebooks, respectively.
As for SSVQ, we have to store the 1-bit sign mask $b_{s}$. The CR is thus calculated as:
$\text{CR}_{\text{SSVQ}}=\frac{O\cdot I\cdot b_{f}}{b_{a}+b_{c}+b_{s}}=\frac{d_{2}\cdot b_{f}}{d_{2}+\lceil\log_{2}(k_{2})\rceil+\frac{k_{2}\cdot d_{2}^{2}\cdot q_{c}}{O\cdot I}}$ (9)
We set different $k$ and $d$ to ensure comparable average bit-width between VQ and SSVQ. Since clustering on all-positive weights is much easier, SSVQ requires much fewer index bits (i.e. smaller $k$ and larger $d$ ). We provide detailed $k\&d$ settings of all experiments in Sec. 12 in the Supplementary Material.